Find the general solution of the following differential equations using method of undetermined coefficients: , (ii) 2y''−5y' = x^2 +5e^−4x,


1
Expert's answer
2022-02-28T17:54:20-0500

Corresponding homogeneous differential equation


"2y''\u22125y' =0"

Characteristic (auxiliary) equation


"2r^2-5r=0"

"r_1=0,r_2=5\/2"

The general solution of the homogeneous differential equation is


"y_h=c_1+c_2e^{5x\/2}"

Find the particular solution of the non homogeneous differential equation


"y_p=Ax^3+Bx^2+Cx+De^{-4x}"

"y_p'=3Ax^2+2Bx+C-4De^{-4x}"

"y_p''=6Ax+2B+16De^{-4x}"

Substitute


"6Ax+2B+16De^{-4x}"

"-15Ax^2-10Bx-5C+20De^{-4x}"

"=x^2+5e^{-4x}"

"-15A=1"

"6A-10B=0"

"2B-5C=0"

"36 D=5"

"A=-1\/15, B=-1\/25, C=-2\/125, D=5\/36"

"y_p=-\dfrac{1}{15}x^3-\dfrac{1}{25}x^2-\dfrac{2}{125}x+\dfrac{5}{36}e^{-4x}"

The general solution of the non homogeneous differential equation is


"y=c_1+c_2e^{5x/2}-\dfrac{1}{15}x^3-\dfrac{1}{25}x^2-\dfrac{2}{125}x+\dfrac{5}{36}e^{-4x}"